OpenStudy (anonymous):

E(6, -2) F (-2, 4) find the distance between the two points

hero (hero):

Once again......same problem. This can be solved a lot quicker with an xy coordinate plane

OpenStudy (anonymous):

oh oops d = ((6 - (-2))^2 + (-2 - (4))^2)^0.5 d = (64 + 36)^0.5 d = 100^0.5 = 10

hero (hero):

I don't even want to know where the .5 exponent comes from

OpenStudy (anonymous):

i know but i like doing it this way better,

hero (hero):

Trust me, if I showed you the way I do it, you would like it.

OpenStudy (anonymous):

well okay

hero (hero):

Wait, let me see if I can use the drawing board to show you.....

OpenStudy (anonymous):

k

hero (hero):